OpenStudy (anonymous):

3x - 4y = 7 3x + 2y = -5 When the second equation is subtracted from the first, the resulting equation is _____. A. -6y = -12 B. -6y = 12 C. -2y = -12 D. 2y = 12

OpenStudy (anonymous):

most days -4-2=-6 and 7+5=12 does that help

OpenStudy (anonymous):

I don';t understand ... @Michiganboy21

OpenStudy (anonymous):

ok look at it like this x\[3-3=0 \] y\[-4-2=-6\] coefficient\[7+5=12\]

OpenStudy (anonymous):

\[3x-4y=7\\3x+2y=-5\] \[3x-3x+2y+4y=-5-7\] \[\boxed{6y=-12\Leftrightarrow-6y=12}\]